OpenStudy (trwatkins1):

What value of m makes the equation true? [–6 + (–13)] + 8 = –6 + (m + 8) m =

OpenStudy (anonymous):

solve left side [-6+(-13)] +8 (-6-13)+8 -19+8 -11 So you have to make the right side =-11 -6 +(m+8)=-11 m+8=-5 m=-13

OpenStudy (mrnood):

do these things in THIS order and post your work here: 1: work out the value of the sum in the square brackets 2: Work out the value of the left hand side of the equation 3: Add 6 to EACH side of the equation 4:Subtract 8 from EACH side of the equation That will give you the value for m

OpenStudy (mrnood):

OR just walk away with m=-13 having SOMEONE ELSE do the work for you